OpenStudy (anonymous):

Given that x+y= 0 is tangent to y=F(x), find F(x) if F'(x) = x

OpenStudy (anonymous):

how do u find the antiderivative of y?

OpenStudy (blockcolder):

Have you taken up the Fundamental Theorem of Calculus?

OpenStudy (campbell_st):

integrate to find F(x) \[\int\limits x dx\] use the tangent properties to evaluate the constant

OpenStudy (anonymous):

yes i have

OpenStudy (dumbcow):

\[F(x) = \frac{1}{2}x^{2} +\frac{1}{2}\]

OpenStudy (anonymous):

how did you get this answer? Can someone explain step by step?

OpenStudy (dumbcow):

\[F(x) = \int\limits_{}^{}x = \frac{1}{2}x^{2} +C\] F'(x) = slope of tangent line = -1 --> x = -1 x+y =0 -1 + y = 0 y = 1 so plug in the point and solve for C 1 = 1/2*(-1)^2 +C